# SplatFlow: Self-Supervised Dynamic Gaussian Splatting in Neural Motion Flow Field for Autonomous Driving

### v1

@@ 3.5 Optical Flow Distillation
$F=\sum_{i=1}^{N}f_{i}\alpha_{i}\prod_{j=1}^{i-1}(1-\alpha_{j}).\vspace{-2.5mm}$ (12)
@@
where $\lambda_{f}$ is a scale factor, and $\mathcal{I}_{next}$ denotes the image at the next time step. $\mathcal{T}$ is used to warp rendered image to next time step according to optical flow. The optical flow $\mathcal{F}$ extracted from foundational model is only required during training and is not needed during inference.
@@ 3.6 Optimization
where the accumulated opacity denotes as $O=\sum_{i=1}^{N}\alpha_{i}\prod_{j=1}^{i-1}(1-\alpha_{j})$.
@@ 4.1 Datasets, Metrics and Baselines
We compare our method with extensive baselines including S-NeRF [21], StreetSurf [5], 3DGS [6], NSG [10], Mars [19], SUDS [17], EmerNerf [24], PVG [1], StreetGaussian [23] on the Waymo and KITTI benchmakrs. More implementation details and results are provided in the supplementary material.
@@ 4.2 Evaluation on Waymo Open Dataset
Following PVG [1], we evaluated our SplatFlow against baselines using the Waymo Open dataset. Table 1 summarizes the average metrics for both image reconstruction and novel view synthesis tasks on selected dynamic scenes. SplatFlow consistently outperforms all baselines across all evaluated metrics for both tasks. In image reconstruction, SplatFlow achieves high image quality with the highest scores: PSNR at 33.64, SSIM at 0.951, and LPIPS at 0.198. For novel view synthesis, SplatFlow produces high-quality renderings of unseen timestamps, reaching PSNR of 28.71, SSIM of 0.874 and LPIPS of 0.239. These improvements are visually validated in Fig. 4 and Fig. 9 (dynamic only), showing exceptional clarity in details for both static and dynamic regions. Compared with baseline methods, which show motion artifacts such as ghosting and blur in regions of dynamic elements, our method preserves high-quality textures and fine details, producing more accurate renderings at novel viewpoints.
@@
w/o NMFF prior 27.28 0.963 0.056
w/o NMFF optimization 27.97 0.975 0.036
w/o optical flow distillation 28.51 0.980 0.028
Full 28.90 0.984 0.026
@@ 4.4 Ablation Study
To demonstrate the effectiveness of each component in SplatFlow, we performed ablation studies on the Waymo dataset. Our main contributions, NMFF prior, NMFF optimization, and optical flow distillation are analyzed in this study to assess their individual impact. We trained four different variations: 1) SplatFlow without pretraining on LiDAR data; 2) SplatFlow without optimization with 4D Gaussians on image data; 3) SplatFlow without optical flow distillation from foundational model; 4) SplatFlow full method. Table 5 presents the average evaluation metrics across test scenes, showing results for both entire scenes and dynamic objects (indicated as *). Ablation study validates the contribution of NMFF prior, NMFF optimization, optical flow distillation to overall performance.
@@ 5 Conclusion
In this supplementary material, we first provide more implementation details of SplatFlow in Appendix 6. Second, we provide more visualization of 3D LiDAR points within NMFF in Appendix 7. Third, we present more detailed comparison visualizations of dynamic object synthesis from novel views in Appendix 8. Fourth, we include visualizations of rendered RGB image, optical flow and depth in Appendix 9. Fifth, runtime performance comparisons are provided in Appendix 10. Finally, video demonstrations are included in Appendix 11.
@@ 7 Visualization of LiDAR Points in NMFF
8 Detailed Visual Comparison
We present more visual comparison details of dynamic object synthesis from novel views, showcasing results on the Waymo dataset in Fig. 11 and on the KITTI dataset in Fig. 12. As shown, SplatFlow generates sharper images with fewer blurred artifacts, particularly for high-speed vehicles, compared to the baselines.
@@ 9 Rendered Depth and Flow Visualization
We also present visualization of rendered RGB image, optical flow, and depth on the Waymo dataset in Fig. 13 and 14, and on the KITTI dataset in Fig. 15. In these Figures, the first row displays the rendered RGB images, the second row shows the rendered optical flow, and the third row presents the rendered depth, all generated by SplatFlow. As shown, our SplatFlow renders sharp, clear, and dense optical flow and depth images in dynamic driving scenarios.
@@ 11 Video Demos
We include five video demos in the supplementary material according to size limitation.
@@
Demos 3 and 4 showcase the rendered images, optical flow, and depth produced by our SplatFlow in dynamic driving scenarios from the Waymo and KITTI datasets. In these videos, the first row shows the G.T RGB images. The second row displays the rendered RGB images, the third row shows the rendered optical flow, and the fourth row presents the rendered depth, all generated by our SplatFlow. For a clearer visualization, video demos 3 and 4 are played at 0.1x speed.

### v2

@@ 3.5 Optical Flow Distillation
$F=\sum_{i=1}^{N}f_{i}\alpha_{i}\prod_{j=1}^{i-1}(1-\alpha_{j}).\vspace{-3mm}$ (12)
@@
where $\lambda_{f}$ is a scale factor, and $\mathcal{I}_{next}$ denotes the image at the next time step. $\mathcal{T}$ is used to warp rendered image to next time step according to optical flow. The optical flow $\mathcal{F}$ extracted from foundation model is only required during training and is not needed during inference.
@@ 3.6 Optimization
where the accumulated opacity denotes as $O=\sum_{i=1}^{N}\alpha_{i}\prod_{j=1}^{i-1}(1-\alpha_{j})$. Note the sky mask is only required during training and is not needed during inference.
@@ 4.1 Datasets, Metrics and Baselines
We compare our method with extensive baselines including S-NeRF [24], StreetSurf [5], 3DGS [7], NSG [12], Mars [22], SUDS [20], EmerNerf [26], PVG [1], StreetGaussian [25] on the Waymo and KITTI benchmarks. See supplementary material for more details and results.
@@ 4.2 Evaluation on Waymo Open Dataset
Following PVG [1], we evaluated our SplatFlow against baselines using the Waymo Open dataset. Table 1 summarizes the average metrics for both image reconstruction and novel view synthesis tasks on selected dynamic scenes. SplatFlow consistently outperforms all baselines across all evaluated metrics for both tasks. In image reconstruction, SplatFlow achieves high image quality with the highest scores: PSNR at 33.64, SSIM at 0.951, and LPIPS at 0.198. For novel view synthesis, SplatFlow produces high-quality renderings of unseen timestamps, reaching PSNR of 28.71, SSIM of 0.874 and LPIPS of 0.239. These improvements are visually validated in Fig. 4 and Fig. 9 (dynamic only), showing exceptional clarity in details for both static and dynamic regions. The dynamic objects in Fig. 9 are selected from a distant background car in an extremely zoomed-in view. Compared with baselines which show artifacts such as ghosting and blur in regions of dynamic elements, our method preserves high-quality textures and fine details, producing more accurate renderings at novel viewpoints.
@@
w/o NMFF prior 27.28 0.827 0.317
w/o NMFF optimization 27.97 0.843 0.254
w/o optical flow distillation 28.51 0.861 0.232
Full 28.90 0.863 0.231
@@ 4.4 Ablation Study
To demonstrate the effectiveness of each component in SplatFlow, we performed ablation studies on the Waymo dataset. Our main contributions, NMFF prior, NMFF optimization, and optical flow distillation are analyzed in this study to assess their individual impact. We trained four different variations: 1) SplatFlow without pretraining on LiDAR data; 2) SplatFlow without optimization with 4D Gaussians on image data; 3) SplatFlow without optical flow distillation from foundational model; 4) SplatFlow full method. Table 5 presents the average evaluation metrics across test scenes used in Table 2, showing results for both entire scenes and dynamic objects (indicated as *). Ablation study validates the contribution of each component to the overall performance.
@@ 5 Conclusion
In this supplementary material, we provide more implementation details of SplatFlow in Appendix 6. We provide more visualization of 3D LiDAR points within NMFF in Appendix 7. We present more detailed comparison visualizations of dynamic object synthesis from novel views in Appendix 8. We provide more visualizations of dynamic objects decomposition in Appendix 9. We present visualizations of the novel view synthesis on newly generated ego-car trajectories in Appendix 10. We include visualizations of rendered RGB image, optical flow and depth in Appendix 11. Runtime performance comparisons are provided in Appendix 12. Video demonstrations are included in Appendix 13.
@@ 7 Visualization of LiDAR Points in NMFF
Figure 11: Dynamic object decomposition results of SplatFlow on Waymo. Row1: Rendered scene, Row2: Corresp. Decomposition
Reference View     Shifted Up 0.5m      Shifted Down 0.5m       Shifted Left 0.5m    Shifted Right 0.5m
Figure 12: Novel view synthesis on newly generated ego-car trajectories on Waymo dataset.
@@
8 Visual Comparison of Dynamic Object Synthesis
We present more visual comparison details of dynamic object synthesis from novel views, showcasing results on the Waymo dataset in Fig. 13 and on the KITTI dataset in Fig. 14. The dynamic objects in these figures are selected from a distant background car in an extremely zoomed-in view. As shown, SplatFlow generates sharper images with fewer blurred artifacts, particularly for high-speed vehicles, compared to the baselines.
@@ 8 Visual Comparison of Dynamic Object Synthesis
9 Visualization of Dynamic Object Decomposition
We provide more dynamic object decomposition visualization of SlatFlow in Fig.11. As shown, SplatFlow clearly separates the dynamic objects from the rendered scenes.
10 Visualization of Novel View Synthesis on newly Ego-car Trajectory
We provide novel view synthesis in a more challenging rendering scenario on the Waymo dataset, using newly generated ego-car trajectories. As shown in Fig. 12, the results are produced from an ego-car trajectory that is shifted 0.5 meters higher, 0.5 meters lower, 0.5 meters to the left, and 0.5 meters to the right of the original trajectory. As demonstrated, SplatFlow can render high-quality novel images for these newly generated ego-car trajectories.
@@ 11 Rendered Depth and Flow Visualization
We also present visualization of rendered RGB image, optical flow, and depth on the Waymo dataset in Fig. 15 and 16, and on the KITTI dataset in Fig. 17. In these Figures, the first row displays the rendered RGB images, the second row shows the rendered optical flow, and the third row presents the rendered depth, all generated by SplatFlow. As shown, our SplatFlow renders sharp, clear, and dense optical flow and depth images in dynamic driving scenarios.
@@ 13 Video Demos
We include five video demos in the supplementary material. See our project page: https://sites.google.com/view/splatflow
@@
Demos 3 and 4 showcase the rendered images, optical flow, and depth by our SplatFlow in dynamic driving scenarios from the Waymo and KITTI datasets. In these videos, the first row shows the G.T RGB images. The second row displays the rendered RGB images, the third row shows the rendered optical flow, and the fourth row presents the rendered depth, all generated by our SplatFlow. For a clearer visualization, video demos 3 and 4 are played at 0.1x speed.